# Prediction of Cellular Malignancy Using Electrical Impedance Signatures and Supervised Machine Learning

### v2

@@ 2.2 Machine Learning Models in Bioelectrical Cancer Classification
Supervised ML algorithms can be trained using datasets compiled from literature and clinical studies to capture diverse patient profiles. In this study, the input features are restricted to key electrical properties -dielectric constant, frequency, conductivity, and characteristic relaxation time constant- as the primary aim is to examine how these properties evolve during the transformation of normal cells to tumorous ones. Understanding these bioimpedance signatures may contribute to future strategies for electrically guided cancer detection or intervention [9],[57],[58],[59],[60]. To improve the model’s robustness, future work will explore the use of simulation tool to generate synthetic data and expand the dataset.
In this work, we evaluate three supervised algorithms – Random Forest (RF), Support Vector Machine (SVM), K- Nearest Neighbor (KNN) – bacause they complement one another in handling non-linear high dimensional bioelectrical features. RF offers inherent bootstrapping and feature importance score and SVM is implemented with a radial-bias -function kernel which is effective with small datasets with decision boundaries.
@@ 2.3 Understanding Random Forest Algorithm
Random Forest (RF) is a supervised ML algorithm that constructs an ensemble of decision trees (DT) to improve prediction accuracy and reduce overfitting [61],[62]. Each decision tree is trained on a random subset of data and selects a random subset of features to make its prediction. In classification tasks, the final output is determined by majority voting among the individual trees, enhancing robustness and generalization of the model. This approach not only enhances prediction accuracy but reduces the risk of overfitting, which is critical when working with biological data that may vary across patients or experimental setups. In the context of this study, RF offers a powerful framework for classifying cells based on their dielectric signatures. Supporting the overarching goal of early, non-invasive cancer detection using bioimpedance characteristics.
@@ 2.4 Understanding Support Vector Machine (SVM)
Another widely used classification model in the biomedical signal analysis is the Support Vector Machine (SVM) [63],[64]. SVM are particularly effective in high dimensional spaces and are known for their robustness in handling small-medium sized datasets with clear margins of separation. By constructing a hyperplane that best divides the data into classes, SVM can classify dielectric properties of tissues. When data is not linearly separable, often in the case with biological signals, SVM employs kernel functions (e.g. radial basis function, polynomial kernel) to project the data into a higher dimensional feature space where linear separation becomes feasible.
@@ 2.5 Understanding K-Nearest Neighbor
K-Nearest Neighbor (KNN) is a non-parametric instance-based supervised machine learning algorithm that classifies data points based on the classes of their nearest neighbors in the feature space [65],[66]. The algorithm relies on a user-defined parameter, k, which determines the number of neighboring points considered when assigning a class label. KNN uses various distance metrics, such as Euclidean distance, Manhattan distance, and Minkowski distance-to compute the similarity between instances [67]. These distance functions are critical in determining the neighborhood structure and can significantly impact classification performance. In the case of breast cancer detection using bioimpedance signatures, KNN can serve as a simple effective baseline model for distinguishing between normal, benign and malignant cells. The algorithm’s strength lies in its simplicity and its adaptability to both linear and nonlinear data distributions without requiring explicit model training. Moreover, it can capture local patterns and is particularly useful in cases where the decision boundary is irregular or data driven. For this study, different values of k will be tested to identify the optimal combination that yields the highest classification accuracy, minimizing false positives or negatives, and improve model reliability. Although KNN can be computationally expensive during inference, especially on large datasets, it remains a useful method for comparative analysis in non-invasive bioimpedance -based cancer classification tasks.
@@ 3 Data Collection and Methodology
For this preliminary study, a sample dataset was compiled from 33 scholarly publications using the keyword “dielectric properties of cancer cells” in academic search engines such as Google Scholar and PubMed Central (PMC). Each manuscript was screened for quantitative data on dielectric properties of normal (healthy), benign and malignant cells. The primary parameters extracted included relative permittivity ($\varepsilon_{r}$), characteristic relaxation time constant ($\tau_{p}$), and conductivity ($\varepsilon$). The frequency range of the datasets spanned from 150 kHz to 20 GHz. Initial analysis revealed that both relative permittivity and conductivity were significantly higher in breast cancer cells compared to their healthy counterparts [2],[9]. This difference highlights the potential of dielectric properties as discrimination biomarkers for cancer classification. These compiled dielectric parameters were then pre-processed for ML model development. Data from different studies were standardized to ensure consistent units and measurement scales. The resulting standardized datasets was subsequently divided into training and testing subsets, forming the foundation for evaluating the classification performance of various supervised learning algorithms, including RF, SVM and KNN. Three supervised ML algorithms were evaluated using the compiled datasets: RF, SVM and KNN. The performance of these algorithms was assessed using four key metrics: accuracy (ACC), F1 score, recall and precision [68],[69],[70],[71]. Accuracy measures the proportion of correct prediction relative to the total prediction made by the model and is generally more reliable for balanced datasets. However, given the limitations and potential imbalance in the literature -derived datasets, recall and precision provide additional insight. Recall is defined as the ratio of correctly classified actual positives to all actual positives, while precision is the ratio of correctly classified actual positives to all predicted positives [72]. The F1-score, calculated as the harmonic means of precision and recall, offers a balanced evaluation of both metrics [73]. As outlined earlier, the primary objective of this study is to compare the predictive performance of the three models, to determine the pathological status of the breast cell, based on the permittivity and conductivity measurement. The algorithm will be used to optimize performance and the corresponding evaluation metrics will be used to identify the model with the highest predictive capability, thereby reducing the risk of false positives and false negatives. For the RF model, the scikit-learn library was used with default parameters [74]. The dataset was partitioned into 80% training and 20% testing subsets. The number of estimators varied from 1 to 500, while the maximum tree depth varied from 1 to 10. The corresponding F1 score, and accuracy were recorded, and the results are discussed in the following section. A similar procedure was applied to the SVM and the linear SVC models, also implemented using the scikit-learn library with default parameter (e.g. regularization parameter C=1). The maximum number of iterations varied between 10, 50, 100, 200, 500, 1000 and 2000. The corresponding evaluation metrics were analyzed to investigate the relationship between hyperparameter variation and performance improvement, particularly with respect to the F1 score.
@@ 4 Results and Discussion
For RF, the number of estimators varied between 1 to 500 with a step size of 10, and the corresponding evaluation metrics were compared. The total duration of the simulation was 29 seconds. The precision, recall, and accuracy stabilized at approximately 86.67%. The RF model was evaluated for different maximum depths, ranging from 1 to 10, while varying the number of estimators at 10 and 500. The maximum accuracy and F1 score were observed at 90% and 88.3% respectively for a maximum depth of 3 and 4 with 100 estimators. Notably, there was no significant gain in accuracy or F1 score when the number of estimators increased from 100 to 500, indicating a saturation effect. On the other hand, the lowest F1 score of 80% was recorded at maximum depth of 8 with only 10 estimators. This highlights two important insights: (i) increasing the number of estimators from 10 to 100 improves stability and performance by reducing variance; (ii) increasing the maximum depth beyond 4 leads to overfitting, as the model begins to capture noise rather than generalizable patterns. For linear SVC, the total computation time was slightly higher at 1.31 seconds. The evaluation metrics were significantly lower, with 53.85%, accuracy, 28.99% precision, and 37.69% F1 score for 2000 iterations. This demonstrates that the linear SVM is not accurate predictive model. The best performance for the SVM model was obtained with a sigmoid kernel and the default regularization parameter (C=1), yielding 76.57% precision, 69.23% accuracy and 64.4% F1 score. In contrast, the worst performance was observed for the RBF kernel at C=15, where accuracy fell to 36.8%. Interestingly, for the polynomial kernel with C=10, the model achieved a high precision of 90% but a poor F1 score of only 37.69%, indicating strong bias towards certain classes. This highlights the trade-off between precision and balanced classification performance in certain hyperparameter settings.
@@
Random Forest (RF) 88.3% Maximum depth = 3 and 4,
Number of estimators = 100
@@ 5 Conclusion and Future Work
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v3

@@ 2.2 Machine Learning Models in Bioelectrical Cancer Classification
Supervised ML algorithms can be trained using datasets compiled from literature and clinical studies to capture diverse patient profiles. In this study, the input features are restricted to key electrical properties -dielectric constant, frequency, conductivity, and characteristic relaxation time constant- as the primary aim is to examine how these properties evolve during the transformation of normal cells to tumorous ones. Understanding these bioimpedance signatures may contribute to future strategies for electrically guided cancer detection or intervention [Hossain2021],[Hossain2024],[Hossain2021b],[Hossain2022],[Hossain2021c]. To improve the model’s robustness, future work will explore the use of simulation tool to generate synthetic data and expand the dataset.
In this work, we evaluate three supervised algorithms – Random Forest (RF), Support Vector Machine (SVM), K- Nearest Neighbor (KNN) – because they complement one another in handling non-linear high dimensional bioelectrical features. RF offers inherent bootstrapping and feature importance score and SVM is implemented with a radial-bias -function kernel which is effective with small datasets with decision boundaries.
@@ 2.3 Understanding Random Forest Algorithm
Random Forest (RF) is a supervised ML algorithm that constructs an ensemble of decision trees (DT) to improve prediction accuracy and reduce overfitting [Salman2024],[Halabaku2024]. Each decision tree is trained on a random subset of data and selects a random subset of features to make its prediction. In classification tasks, the final output is determined by majority voting among the individual trees, enhancing robustness and generalization of the model. This approach not only enhances prediction accuracy but reduces the risk of overfitting, which is critical when working with biological data that may vary across patients or experimental setups. In the context of this study, RF offers a powerful framework for classifying cells based on their dielectric signatures. Supporting the overarching goal of early, non-invasive cancer detection using bioimpedance characteristics.
@@ 2.4 Understanding Support Vector Machine (SVM)
Another widely used classification model in the biomedical signal analysis is the Support Vector Machine (SVM) [Pisner2020],[Meyer2001]. SVM are particularly effective in high dimensional spaces and are known for their robustness in handling small-medium sized datasets with clear margins of separation. By constructing a hyperplane that best divides the data into classes, SVM can classify dielectric properties of tissues. When data is not linearly separable, often in the case with biological signals, SVM employs kernel functions (e.g. radial basis function, polynomial kernel) to project the data into a higher dimensional feature space where linear separation becomes feasible.
@@ 2.5 Understanding K-Nearest Neighbor
K-Nearest Neighbor (KNN) is a non-parametric instance-based supervised machine learning algorithm that classifies data points based on the classes of their nearest neighbors in the feature space [Bezdek1986],[Zhang2016]. The algorithm relies on a user-defined parameter, k, which determines the number of neighboring points considered when assigning a class label. KNN uses various distance metrics, such as Euclidean distance, Manhattan distance, and Minkowski distance-to compute the similarity between instances [Iqbal2020]. These distance functions are critical in determining the neighborhood structure and can significantly impact classification performance. In the case of breast cancer detection using bioimpedance signatures, KNN can serve as a simple effective baseline model for distinguishing between normal, benign and malignant cells. The algorithm’s strength lies in its simplicity and its adaptability to both linear and nonlinear data distributions without requiring explicit model training. Moreover, it can capture local patterns and is particularly useful in cases where the decision boundary is irregular or data driven. For this study, different values of k will be tested to identify the optimal combination that yields the highest classification accuracy, minimizing false positives or negatives, and improve model reliability. Although KNN can be computationally expensive during inference, especially on large datasets, it remains a useful method for comparative analysis in non-invasive bioimpedance -based cancer classification tasks.
@@ 3 Data Collection and Methodology
For this preliminary study, a sample dataset was compiled from 20 scholarly publications with 535 datasets using the keyword “dielectric properties of cancer cells” in academic search engines such as Google Scholar and PubMed Central (PMC). Each manuscript was screened for quantitative data on dielectric properties of normal (healthy), benign and malignant cells. The primary parameters extracted included relative permittivity ($\varepsilon_{r}$), characteristic relaxation time constant ($\tau_{p}$), and conductivity ($\varepsilon$). The frequency range of the datasets spanned from 150 kHz to 20 GHz. Initial analysis revealed that both relative permittivity and conductivity were significantly higher in breast cancer cells compared to their healthy counterparts [Hossain2020],[Hossain2021]. This difference highlights the potential of dielectric properties as discrimination biomarkers for cancer classification. These compiled dielectric parameters were then pre-processed for ML model development. Data from different studies were standardized to ensure consistent units and measurement scales. The resulting standardized datasets was subsequently divided into training and testing subsets, forming the foundation for evaluating the classification performance of various supervised learning algorithms, including RF, SVM and KNN. Three supervised ML algorithms were evaluated using the compiled datasets: RF, SVM and KNN. The performance of these algorithms was assessed using four key metrics: accuracy (ACC), F1 score, recall and precision [Asif2021],[Hossain2025],[Hossain2025b],[hossain2025c]. Accuracy measures the proportion of correct prediction relative to the total prediction made by the model and is generally more reliable for balanced datasets. However, given the limitations and potential imbalance in the literature -derived datasets, recall and precision provide additional insight. Recall is defined as the ratio of correctly classified actual positives to all actual positives, while precision is the ratio of correctly classified actual positives to all predicted positives [Flach2015]. The F1-score, calculated as the harmonic means of precision and recall, offers a balanced evaluation of both metrics [Diallo2024]. As outlined earlier, the primary objective of this study is to compare the predictive performance of the three models, to determine the pathological status of the breast cell, based on the permittivity and conductivity measurement. The algorithm will be used to optimize performance and the corresponding evaluation metrics will be used to identify the model with the highest predictive capability, thereby reducing the risk of false positives and false negatives. For the RF model, the scikit-learn library was used with default parameters [Pedregosa2011]. The dataset was partitioned into 80% training and 20% testing subsets. The number of estimators varied from 1 to 500, while the maximum tree depth varied from 1 to 10. The corresponding F1 score, and accuracy were recorded, and the results are discussed in the following section. A similar procedure was applied to the SVM and the linear SVC models, also implemented using the scikit-learn library with default parameter (e.g. regularization parameter C=1). The maximum number of iterations varied between 10, 50, 100, 200, 500, 1000 and 2000. The corresponding evaluation metrics were analyzed to investigate the relationship between hyperparameter variation and performance improvement, particularly with respect to the F1 score.
@@ 4 Results and Discussion
The Random Forest (RF) classifier was evaluated by varying the number of estimators from 100 to 500 in increments of 100 as shown in Table 3. The corresponding performance metrics were computed for each configuration. Across the evaluated configurations, the precision recall, and overall classification accuracy converged to approximately 86-87% indicating stable ensemble performance. To further investigate model complexity, the maximum tree depth varied from 1 to 10 while considering estimator values of 100 and 500. The highest classification accuracy ($\approx$87%) and macro averaged F1 score were achieved at higher tree depths ($\geq$ 10), irrespective of the number of estimators. Increasing the maximum depth beyond 10 did not yield a statistically significant improvement in performance, indicating a saturation effect in ensemble learning capacity. In contrast, shallow tree (maximum depth=1) resulted in reduced performance, with a F1-score of approximately 66%, demonstrating underfitting due to limited model representation capability. Furthermore, no substantial performance improvement was observed for depths exceeding 10, suggesting that model complexity saturation occurs beyond this threshold.
Figure 5: Comparative performance analysis of the Random Forest classifier for varying hyperparameter configurations. (a), (d), and (g) report class wise quantitative metrics; (b), (e), and (h) depict training versus cross-validation accuracy trends; and (c), (f), and (i) show confusion matrices. Case I : depth=5, estimators =100; Case II : depth=3, estimators =100; Case III: depth=15, estimators =500.
Fig. 5 (a) –(i) presents the quantitative evaluation results, including (i) class-wise and macro-averaged performance metrics, (ii) the corresponding confusion matrices, and (iii) training and cross validation learning curves for three representative hyperparameter configurations: Case I : Maximum depth=5, Number of estimators =100; Case II: Maximum depth=3, Number of estimators=100, and Case III : Maximum depth=15, number of estimators =500. In all three cases, the training accuracy consistently exceeds the cross-validation accuracy, indicating the presence of moderate overfitting. However, the performance gap decreases for deeper trees, suggesting improved generalization with increased model capacity. A comparative assessment of Fig. 5 (a), (d) and (g) reveals that the RF classifier demonstrates strong discriminative capacity for malignant cells (Class III), achieving F1 scores exceeding 0.88 in Case III. Conversely, the classification performance for benign cells (Class II) remains comparatively lower, with F1 scores of approximately 0.59 and 0.54 for Case I and II, respectively and improving to approximately 0.71 in Case III. The relatively lower benign -class sensitivity highlights the inherent difficulty in distinguishing benign from malignant cellular dielectric signatures. Enhancing benign-class discrimination is clinically significant as improved early stage differentiation may facilitate timely diagnosis and optimized therapeutic interventions.
Table 3: Classification Performance (F1 Score) for Various Tree Depths and Number of Estimators
Estimators Depth = 1 Depth = 5 Depth = 10
100 66% 76% 86%
200 66% 78% 86%
300 70% 76% 87%
400 69% 77% 87%
500 69% 76% 86%
@@
For linear SVC, the total computation time was slightly higher at 1.31 seconds. The evaluation metrics were significantly lower, with 53.85%, accuracy, 28.99% precision, and 37.69% F1 score for 2000 iterations. This demonstrates that the linear SVM is not accurate predictive model. The best performance for the SVM model was obtained with a sigmoid kernel and the default regularization parameter (C=1), yielding 76.57% precision, 69.23% accuracy and 64.4% F1 score. In contrast, the worst performance was observed for the RBF kernel at C=15, where accuracy fell to 36.8%. Interestingly, for the polynomial kernel with C=10, the model achieved a high precision of 90% but a poor F1 score of only 37.69%, indicating strong bias towards certain classes. This highlights the trade-off between precision and balanced classification performance in certain hyperparameter settings.
@@
Random Forest (RF) 87% Maximum depth = 10,
Number of estimators = 400
@@ 5 Conclusion and Future Work
Shadeeb Hossain is the sole proprietor of Shadeeb Engineering Lab, which may have a commercial interest in technologies related to this research.